OpenStudy (anonymous):

Find the solution to the system by the substitution method. Check your answers. 10x + 3y = 8 2x + y =2

OpenStudy (anonymous):

10x+3y=8 10x+5y=4 -2y=4 y=-2 2x-2=2 x=2

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

oops, sorry, should be 10x+3y=8 10x+5y=10 -2y=-2 y=1 2x+1=2 x=1/2